# Virtual Reflections on a Dynamic 2D Eye Model Improve Spatial Reference Identification

### v3

@@ 2.2.4 Participants
The experiment involved 30 participants (4 females and 26 males; aged 21 to 56 years who participated in groups of three; mean age: $32.8\pm 8.52$ years). All participants signed consent forms, and the experiment was reviewed and approved by the Bioethics Committee in Honda’s R&D (100HM-064H). One participant was excluded from accuracy and reaction time analysis due to a lack of recorded button presses.
2.2.5 Setup
@@
Figure 4 shows the schematic of the experimental setup. The experiments were conducted in an office environment in which three participants stood in front of a 65-inch display (Bravia KJ-65X80J; Sony, Japan). Participants stood approximately 2 meters from the display, with a 1-meter distance between one another (Fig. 4(a)). A camera (Azure Kinect DK; Microsoft, USA) was mounted directly above the center of the screen to capture the scene containing the participants, i.e., the eyes’ attention targets. Each participant was equipped with a key button (FS1-P; Edikun, Japan) on their hand, connected to the experimental computer via USB. A computer (Magnus One ECM73070C; Zotac, Hong Kong) managed both the visual stimulus presentation and data collection. The experimental stimulus, an image of a generic robot head equipped with the dynamic eye model of the respective condition, was displayed at the center of the screen on a gray background with a height of 26 cm. Each eye was represented as a 9 × 9 cm square. The diameter of the outer circle of the pupil in the eye was 6.4 cm, and the diameter of the inner circle was 3.5 cm. The display measured 80.6 cm in height and 143.1 cm in width and was positioned on a stand that was about 99.5 cm tall. This positioning ensured that the virtual eye models were at a height of approximately 1.5 meters. This height was chosen to ensure that the participants were in a comfortable position relative to the virtual eye gaze while standing (Figure 4(b)).
(a) Schematic of participants’ positions
(b) Experiment stimuli presented on a TV display
Figure 4: Schematic of the experimental setup: (a): The diagram shows the distance of the three participants from the display and their respective positions. The participants had the freedom to move their bodies or head while in their positions. (b): The illustration depicts the size and height of the visual stimuli presented as cues on the gray display.
2.2.6 Software
@@
We implemented the experiment using the Python programming language [27]. The subsequent data analysis was conducted in Python using the libraries Pandas  [30] for data structuring, Matplotlib [31] and seaborn [32] for visualizations, and statsmodels [33] and scikit-posthocs [34] for statistical analysis.
@@ 3 Results
3.1 Reference identification performance
To test H1 and H2, we used the reference identification accuracy and reaction times as dependent measures. Both are derived entirely from the participants’ button press behavior. The reference identification accuracy for each participant is defined as binary classification accuracy, i.e., as the ratio of the sum of true positive ($TP$) and true negative ($TN$) identifications over the sum of all trials involving that participant.
$\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$ (5)
Therefore, the accuracy does not only take the true positive identification of a reference into account but is also affected by the participants’ ability to identify when they are not being attended. To provide opportunities for the identification of such non-references, we programmed the eye model to "make mistakes" in a subset of trials by not attending to any particular participant but instead focusing on areas located between participants. For such trials, the lack of a button press within the time limit would then count as a true negative identification. The rate of such machine mistakes was set to 10% for each block in part 1 and to 20% for part 2.
We chose to only use a single response button per participant to avoid any coordination-dependent confounders. This button indicated a positive response. Negative responses were given implicitly by not responding within the 3-second time limit. This choice means that reaction times were only available for true positive and false positive responses as indicated by Hypothesis 2.
3.1.1 Accuracy
Table 1: ANOVA table for accuracy data.
@@ 3.1.1 Accuracy
(a) Accuracy in single condition blocks (part 1)
(b) Accuracy in the mixed condition block (part 2)
Figure 5: Box plots showing the distributions of accuracy scores across conditions for single condition blocks (left) and mixed condition blocks (right).
Figure 5 shows the distribution of accuracy values across conditions in both parts of the experiment. The accuracy was high across all conditions for the single condition blocks (Eye only: $\mu$=0.82, sd=0.13; Mirror only: $\mu$=0.91, sd=0.12; Mirror eye: $\mu$=0.94, sd=0.73). Also in the mixed condition block the accuracy values were high across conditions (Eye only: $\mu$=0.81, sd=0.2; Mirror only: $\mu$=0.92, sd=0.15; Mirror eye: $\mu$=0.94, sd=0.094). A two-way ANOVA revealed a significant main effect for display condition (F(2,168)=14.99, p<0.001) but no effect for block type (F(1,168)=0.05, p=0.82) or interaction between display condition and block type (F(2,168)=0.09, p=0.91). Post-hoc comparisons revealed a significant difference between the eye-only and mirror-only (p < 0.01) and between the eye-only and the mirror eye condition (p < 0.001).
These results support our first hypothesis and additionally indicate an accuracy advantage of a dynamic mirroring mode over pure eye-based referencing. There are no indications of an influence of cue-type predictability on the average accuracy.
3.1.2 Reaction times
Table 2: ANOVA table for reaction time data.
@@ 3.1.2 Reaction times
(a) Reaction times in single condition blocks (part 1)
(b) Reaction times in the mixed condition block (part 2)
Figure 6: Box plots showing the distributions of reaction times across conditions for single condition blocks (left) and mixed condition blocks (right).
Figure 6 shows the distribution of reaction times across conditions in experiments 1 and 2. Average reaction times were similar across all conditions for the single condition blocks of part 1 (Eye only: $\mu$=1.29s, sd=0.29; Mirror only: $\mu$=1.38, sd=0.2; Mirror eye: $\mu$=1.34, sd=0.24). Reaction times in the mixed condition block of part 2 were comparable to those of the first part, especially for the two mirror conditions (Eye only: $\mu$=1.46s, sd=0.32; Mirror only: $\mu$=1.4, sd=0.32; Mirror eye: $\mu$=1.35, sd=0.27). A two-way ANOVA did not reveal any significant effects for display condition (F(2,168)=0.35, p=0.7), block type (F(1,168)=2.45, p=0.12) or interaction between display condition and block type (F(2,168)=1.59, p=0.2).
These results are not indicative of any reaction time disadvantage of the mirror-only and mirror-eye modes compared to the eye-only mode and hence do not support the second hypothesis. There are no clear indications of an influence of cue-type predictability on the average reaction time.
@@ 3.2 User experience
In this experiment, the 8-item version of the User Experience Questionnaire (UEQ-S; see Table 3 [35]) was used to evaluate participants’ experience with the virtual eyes. After completing all experimental blocks, participants filled out the UEQ-S, which consists of eight items, four representing a pragmatic scale and the other four representing a hedonic scale. Each item is rated on a 7-point semantic differential scale, ranging from -3 (most negative evaluation) to +3 (most positive evaluation).
Table 3: The items in the UEQ-S
Negative term 7-point scale Positive term
obstructive o o o o o o o supportive
complicated o o o o o o o easy
inefficient o o o o o o o efficient
clear o o o o o o o confusing
boring o o o o o o o exciting
not interesting o o o o o o o interesting
conventional o o o o o o o inventive
usual o o o o o o o leading edge
(a) Pragmatic quality
(b) Hedonic quality
Figure 7: Box plots for UEQ-S rating of pragmatic and hedonic quality across the three different display conditions of the eye models. Different colors represent evaluation ranges: values between -0.8 and 0.8 indicate neutral ratings (grey), values above 0.8 signify positive ratings (green), and values below -0.8 indicate negative ratings (red).
Figures 7 and 8 present the results of the UEQ-S ratings by 30 participants who evaluated the user experience across three different display conditions of the eye models. Figures 7(a) and 7(b) illustrate the pragmatic and hedonic quality, respectively. In the box plot figures, different background colors indicate varying evaluation ranges. Values between -0.8 and 0.8 (grey background) indicate a neutral evaluation on the corresponding scale. Values greater than 0.8 (green background) signify a positive evaluation, while values below -0.8 (red background) indicate a negative evaluation. Table 4 shows the scale means for the UEQ-S ratings by 30 participants across three different display conditions of eye models. Standard deviations are shown in parentheses. Bold font indicates a statistically significant ($p<0.001$) difference from other conditions.
Table 4: Scale scores means and standard deviations of the UEQ-S (N = 30)
@@
Eye-only 0.67 (1.27) -0.26 (1.34) 0.20 (1.12)
Mirror-only 0.71 (1.59) -0.22 (1.46) 0.25 (1.12)
Mirror-eye 1.29 (1.38) 1.30 (0.92) 1.30 (0.98)
For the eye-only condition, the mean score was 0.67 for pragmatic quality and -0.26 for hedonic quality. These scores are similar to those in the mirror-only condition for which the means were 0.71 for pragmatic quality and -0.22 for hedonic quality. In contrast, the mirror-eye condition yielded the highest rating among the three conditions, with mean scores of 1.29 for pragmatic quality and 1.3 for hedonic quality. A one-way ANOVA revealed a significant difference in hedonic quality across the three conditions ($F(2,87)=14.86,p<0.001$). Post-hoc multiple comparisons showed a significant difference between the mirror-eye and the other two conditions (both $p<0.001$). For pragmatic quality, while the mirror-eye condition had a higher mean (1.29, indicating a positive rating in the green zone), no significant differences were found between the three conditions ($F(2,87)=1.81,p=0.17$).
Figure 8: Box plot showing the distribution of the overall measure of user experience across the three different display conditions of the eye models. Values between -0.8 and 0.8 indicate neutral ratings (grey), values above 0.8 signify positive ratings (green), and values below -0.8 indicate negative ratings (red).
Figure 8 displays the overall user experience evaluation. The mirror-eye condition received a mean score of 1.30, while the eye-only and mirror-only conditions scored 0.20 and 0.25, respectively. ANOVA results indicated a significant difference in overall user experience across the three conditions ($F(2,87)=9.93,p<0.001$). Post-hoc multiple comparisons confirmed that the mirror-eye condition differed significantly from the other two conditions (both $p<0.001$). Additionally, the internal consistency of the pragmatic quality and hedonic quality scales for the mirror-eye condition was sufficiently high, with Cronbach Alpha values of 0.88 (pragmatic quality) and 0.81 (hedonic quality).
In sum, the results from the UEQ-S ratings indicate a more positive user experience for the mirror eye condition than for the other two conditions and thus support our third hypothesis.
@@ 4 Discussion
Here we introduced an approach for overcoming spatial referencing challenges of 2D eye models as communication tools in human-machine interaction. By augmenting a screen-based eye model with a dynamic reflection-like overlay that is contingent on eye movements, pupil-based pointing is complemented by direct scene context to improve spatial reference identification for observers. We conducted a user study with 30 participants who were asked to carry out a group interaction task that relied on an identification of spatial references produced by screen-based eye models to compare this "mirror eye" model to a pure eye model without mirroring capabilities and a pure dynamic mirror model without eye-like features.
@@ 4.1 Reference identification accuracy and speed
Although accuracy was high for all tested eye models, the two eye models with mirror-like features led to significantly higher reference identification accuracy than a pure non-reflective eye model. Despite the added contextual information, the observers of augmented eye models needed to process, the study did not reveal any significant impact of mirroring on reaction times. The lack of noticeable impact on reaction times stands in contrast to our second hypothesis. One possible explanation lies in the reference target class used in our study: The eye models always focused on faces.
Human brains have evolved highly dedicated face processing abilities [36], likely due to their social role, allowing us to recognize target faces after as little as  260ms [37]. Furthermore, the seen gaze direction has been found to modulate the activity of regions associated with face processing [38], which could suggest that a linkage of the two, as introduced here, might be exploiting fast visual processing networks. In addition, occasionally seeing their own faces might have facilitated the processing of other participants’ faces [39], which could have reduced the risk of false positive identifications.
However, the potential links to neural face processing and associated findings also highlight one limitation of the present study: Because faces were the only spatial reference targets, we gain no insights about the utility of the approach for conveying references to other object categories that are likely to have fewer dedicated neural processing resources.
@@ 4.2 User experience
Our findings regarding accuracy and response speed are consistent with user experience ratings obtained via the UEQ-S questionnaire.
The UEQ-S results indicate that the mirror-eye condition was attributed higher pragmatic and hedonic qualities than the other conditions, supplying the means for the step from a neutral to a positive user experience. Overall, the statistical analysis revealed that the user experience with the mirror-eye condition was significantly better than with the eye-only or mirror-only conditions. Participants hence appear to have largely appreciated the performance improvements accompanied by the mirror augmentation.
Notably, one participant differed in this regard by rating both mirror conditions lower than the pure eye model. This participant explained afterwards, that seeing oneself in the eyes evoked a strong sense of being observed that made them feel uncomfortable. Although this response came from a 1/30 minority of the participants, the potential for such responses may require additional consideration in future interaction designs involving similar features.
@@ 4.3 Mental state transparency
Differential emotional responses to seeing oneself in the mirror have been described before [40, 41, 42] and could account for some variability in user experience. Another factor that may contribute to a negative sense of being observed is an attribution of judgment or malevolence to the observing entity. Because we were mainly interested in the mirror eyes’ utility for improving spatial ocular referencing we tried to keep the expression of the eye model itself neutral. However, we cannot rule out that individual users may attribute diverse mental states to the agent driving an eye model. This highlights further potential in the utilization of augmented eye models such as the mirror eyes.
Eyes are often labeled as windows to the mind. Here we only used them to convey spatial attention but human ocular behavior can also reflect other mental states. A transfer of ocular behavior that conveys specific mental states to eye models, e.g., to counteract false associations, has already been demonstrated. For example, carefully crafted eye-gazing behaviors, when integrated with external information in avatars or virtual characters, have been used in virtual reality settings to subtly convey cognitive states and enhance social presence [43, 44, 45]. Moreover, when paired with nonverbal cues like gestures and facial expressions, gaze inference models show great promise in human-computer interaction applications [46, 47]. These models also have the potential to express an agent’s personality traits [46] and emotional states, such as fear and anxiety [48].

### v4

@@ 3.6 Participants
The experiment involved 30 participants (4 females and 26 males; aged 21 to 56 years; mean age: $32.8\pm 8.52$ years) with normal or corrected-to-normal vision, who participated in groups of three. Participants had no prior exposure to the experimental setup or eye model. All participants signed consent forms, and the experiment was reviewed and approved by the Bioethics Committee in Honda’s R&D (100HM-064H). One participant was excluded from accuracy and reaction time analysis due to a lack of recorded button presses.
3.7 Setup
@@
Figure 6 shows the schematic of the experimental setup. The experiments were conducted in an office environment in which three participants stood in front of a 65-inch display (Bravia KJ-65X80J; Sony, Japan). Participants stood approximately 2 meters from the display, with a 1-meter distance between one another (Fig. 6b). The positions were selected to provide a comfortable social distance while still keeping participants close enough to each other and far enough from the display to create some gaze direction ambiguity. A camera (Azure Kinect DK; Microsoft, USA) was mounted directly above the center of the screen to capture the scene containing the participants, i.e., the eyes’ attention targets. Each participant was equipped with a key button (FS1-P; Edikun, Japan) on their hand, connected to the experimental computer via USB. A computer (Magnus One ECM73070C; Zotac, Hong Kong) managed both the visual stimulus presentation and data collection. The experimental stimulus, an image of a generic robot head equipped with the dynamic eye model of the respective condition, was displayed at the center of the screen on a gray background with a height of 26 cm (Figure 6a). Each eye was represented in a 9 × 9 cm square. The diameter of the sclera of each eye was 8.8 cm, the diameter of the iris (outer circle) was 6.4 cm, and the diameter of the pupil (inner circle) was 3.5 cm. The reflection opacity in the Mirrors-Only and Mirror Eyes  conditions was set to 60%. The display measured 80.6 cm in height and 143.1 cm in width and was positioned on a stand that was about 99.5 cm tall. This positioning ensured that the eye models were at a height of approximately 1.5 meters, which allowed participants to observe them comfortably from a standing position.
3.8 Software
@@
We implemented the experiment using the Python programming language [31]. The subsequent data analysis was conducted in Python using the libraries Pandas  [38] for data structuring, Matplotlib [39] and seaborn [40] for visualizations, and statsmodels [41], scikit-posthocs [42], and pingouin [43] for statistical analysis.
@@ 4 Results
Table 1: Summary of Two-Way ANOVA results for accuracy values.
@@
(a) Single condition blocks (part 1)
(b) Mixed condition block (part 2)
Figure 7: Box plots showing the distributions of accuracy scores across conditions for single (left) and mixed condition blocks (right).
Table 2: Summary of Two-Way ANOVA results for reaction times.
@@
(a) Single condition blocks (part 1)
(b) Mixed condition block (part 2)
Figure 8: Box plots showing the distributions of reaction times across conditions for single (left) blocks and the mixed condition block (right).
4.1 Reference identification performance
4.1.1 Accuracy
Figure 7 shows the distribution of accuracy values across conditions in both parts of the experiment. The accuracy was high across all conditions for the single condition blocks (Eyes-Only: $\mu=0.82,sd=0.13$; Mirrors-Only: $\mu=0.91,sd=0.12$; Mirror Eyes: $\mu=0.94,sd=0.07$). Also in the mixed condition block the accuracy values were high across conditions (Eyes-Only: $\mu=0.81,sd=0.20$; Mirrors-Only: $\mu=0.92,sd=0.15$; Mirror Eyes: $\mu=0.94,sd=0.09$). A two-way ANOVA (see Table 1) revealed a significant main effect for display condition ($F(2,168)=14.99,p<0.001$) but no effect for block type ($F(1,168)=0.05,p=0.82$), or interaction between display condition and block type ($F(2,168)=0.09,p=0.91$). Post-hoc comparisons with Holm-Bonferroni adjusted p-values revealed a significant difference between the Eyes-Only and Mirrors-Only ($p<0.01$) and between the Eyes-Only and the Mirror Eyes condition ($p<0.001$). These results support our first hypothesis and additionally indicate an accuracy advantage of a dynamic mirroring mode over pure eye-based referencing. There are no indications of an influence of cue type predictability on the average accuracy.
4.1.2 Reaction times
Figure 8 shows the distribution of reaction times across conditions in experiment parts 1 and 2. Average reaction times were similar across all conditions for the single condition blocks of experiment part 1 (Eyes-Only: $\mu=1.29,sd=0.29$; Mirrors-Only: $\mu=1.38,sd=0.20$; Mirror Eyes: $\mu=1.34,sd=0.24$). Reaction times in the mixed condition block of experiment part 2 were comparable to those of the first part, especially for the two mirror conditions (Eyes-Only: $\mu=1.46,sd=0.32$; Mirrors-Only: $\mu=1.40,sd=0.32$; Mirror Eyes: $\mu=1.35,sd=0.27$). A two-way ANOVA (see Table 2) did not reveal any significant effects for display condition ($F(2,168)=0.35,p=0.70$), block type ($F(1,168)=2.45,p=0.12$), or interaction between display condition and block type ($F(2,168)=1.59,p=0.20$).
These results are not indicative of any reaction time disadvantage of the Mirrors-Only and Mirror Eyes modes compared to the Eyes-Only mode and hence do not support the second hypothesis. There are no clear indications of an influence of cue type predictability on the average reaction time.
@@ 4.2 User experience
Table 3: Score means and standard deviations of the UEQ-S (N = 30)
@@
Eyes-Only 0.67 (1.27) -0.26 (1.34) 0.20 (1.12)
Mirrors-Only 0.71 (1.59) -0.22 (1.46) 0.25 (1.12)
Mirror Eyes 1.29 (1.38) 1.30 (0.92) 1.30 (0.98)
Bold values denote statistically significant differences ($p<0.01$).
Figure 9: Box plots for UEQ-S rating of pragmatic and hedonic quality, and overall user experience across the three different conditions of the eye models. Different background colors represent evaluation ranges: values between 0.8-0.8 and $0.8$ indicate neutral ratings (grey), values above $0.8$ signify positive ratings (green), and values below 0.8-0.8 indicate negative ratings (red).
Figure 9 presents the results of the UEQ-S ratings of 30 participants who evaluated the user experience across three different display conditions of the eye models, which were referred to as ”Eyes”, ”Mirrors”, and ”Mirror Eyes” during the study. Table 3 shows the scale means for the UEQ-S ratings. Standard deviations are shown in parentheses.
For the Eyes-Only condition, the mean score was 0.67 for pragmatic quality and -0.26 for hedonic quality. These scores are similar to those in the Mirrors-Only condition for which the means were 0.71 for pragmatic quality and -0.22 for hedonic quality. In contrast, the Mirror Eyes condition yielded the highest rating among the three conditions, with mean scores of 1.29 for pragmatic quality and 1.3 for hedonic quality. A one-way ANOVA revealed a significant difference in hedonic quality across the three conditions ($F(2,87)=14.86,p<0.001$). Post-hoc multiple comparisons with Holm-Bonferroni adjusted p-values showed a significant difference between the Mirror Eyes and the other two conditions (both $p<0.001$). For pragmatic quality, while the Mirror Eyes condition had a higher mean (1.29, indicating a positive rating in the green zone), no significant differences were found between the three conditions ($F(2,87)=1.81,p=0.17$).
For the overall user experience evaluation, the Mirror Eyes condition received a mean score of 1.30, while the Eyes-Only and Mirrors-Only conditions scored 0.20 and 0.25, respectively. ANOVA results indicated a significant difference in overall user experience across the three conditions ($F(2,87)=9.93,p<0.001$). Post-hoc multiple comparisons with Holm-Bonferroni adjusted p-values confirmed that the Mirror Eyes condition differed significantly from the other two conditions (both $p<0.001$). The internal consistency of the pragmatic quality and hedonic quality scales for the Mirror Eyes condition was sufficiently high, with Cronbach Alpha values of 0.88 (pragmatic quality) and 0.81 (hedonic quality).
In sum, the results of the UEQ-S ratings indicate a more positive user experience for the Mirror Eyes condition than for the other two conditions and thus support our third hypothesis.
@@ 5 Discussion
Here, we introduced an approach for overcoming spatial referencing challenges of 2D eye models as communication tools in human-machine interaction. By augmenting a screen-based eye model with a dynamic reflection-like overlay that is contingent on eye movements, pupil-based pointing is complemented by direct scene context to improve spatial reference identification for observers. We conducted a user study with 30 participants who were asked to carry out a group interaction task that relied on an identification of spatial references produced by screen-based eye models to compare this ”Mirror Eyes” model to a pure eye model without mirroring capabilities and a pure dynamic mirror model without eye-like features.
@@ 5.1 Reference identification accuracy and speed
Although accuracy was high for all tested eye models, the two eye models with mirror-like features led to significantly higher reference identification accuracy than a pure non-reflective eye model. Despite the added contextual information, the observers of augmented eye models needed to process, the study did not reveal any significant impact of mirroring on reaction times. The lack of noticeable impact on reaction times stands in contrast to our second hypothesis, which assumed a speed reduction for reference identification as an expression of a potential trade-off for the additional information that needs to be processed. The absence of a mirroring-related speed trade-off might in part be explained by the practice trials prior to the experiment. These could have given participants an opportunity to identify and focus on the most informative features. Future work should facilitate the separation of dependent measures from early and later trials of each experimental condition to identify potential learning effects.
Another possible explanation for the lack of an apparent trade-off lies in the reference target class used in our study: The eye models always focused on faces. Human brains have evolved highly dedicated face processing abilities [44], likely due to their social role, allowing us to recognize target faces after as little as 260 ms [45]. Furthermore, the seen gaze direction has been found to modulate the activity of regions associated with face processing [46], which could suggest that a linkage between the two, as introduced here, might be exploiting fast visual processing networks. In addition, occasionally seeing their own faces might have facilitated the processing of other participants’ faces [47], which could have reduced the risk of false positive identifications.
However, the potential links to neural face processing and associated findings also highlight one limitation of the present study: Because faces were the only spatial reference targets, we gain no insights about the utility of the approach for conveying references to other object categories that are likely to have fewer dedicated neural processing resources. Future investigations should therefore expand the set of gaze target object categories.
@@ 5.2 User experience
Our findings regarding accuracy and response speed are consistent with user experience ratings obtained via the UEQ-S questionnaire. The UEQ-S results indicate that the Mirror Eyes condition was attributed significantly higher hedonic qualities than the other conditions, supplying the means for the step from a neutral to a positive user experience. Overall, the statistical analysis revealed that the user experience with the Mirror Eyes condition was significantly better than with the Eyes-Only or Mirrors-Only conditions. The participants hence appear to have largely appreciated the support given by the Mirror Eyes condition.
Notably, a single participant differed in this regard by rating both mirror conditions lower than the pure eye model. This participant explained afterwards, that seeing oneself in the eyes evoked a strong sense of being observed that made them feel uncomfortable. Although this response came from a 1/30 minority of the participants, the potential for such responses may require additional consideration in future interaction designs involving similar features. Furthermore, the male-skewed participant sample may have constrained the range of experiences and behaviors represented in our findings.
@@ 5.3 Mental state transparency
Differential emotional responses to seeing oneself in the mirror have been described before [48, 49, 50] and could account for some variability in user experience. Another factor that may contribute to an adverse sense of being observed is an attribution of judgment or malevolence to the observing entity. Because we were mainly interested in the Mirror Eyes’ utility for improving spatial ocular referencing, we kept the expression of the eye model neutral. However, we cannot rule out the possibility that individual users may attribute diverse mental states or personalities [51] to the agent driving an eye model. Close associations between eyes and mental states may impair some interaction scenarios, but also highlight further potential in the utilization of augmented eye models such as the Mirror Eyes.
Often described as windows to the mind, the eyes can reflect a larger set of mental states than just spatial attention. For instance, carefully designed gaze patterns, especially when integrated with contextual information in avatars or virtual agents, have been used in VR to subtly convey cognitive states and enhance social presence [52, 53]. Moreover, when combined with non-verbal cues such as gestures and facial expressions, gaze inference models show great promise in human-computer interaction applications [54, 55], and can also be used to express personality traits [54] or emotional states like fear and anxiety [56]. Nevertheless, close associations between eyes and mental states may also limit possible applications.
5.4 Cue saliency and attention
Eyes by themselves are strong attention attractors [3, 4, 5]. Reflection-like overlays may increase the salience of eye models even further, especially for reflections of other strong attractors like faces [57, 58]. The salience of any machine expression deserves consideration in interaction design because it affects the balance between observability and distraction. Typically, a high expression salience makes an observer more likely to look at a face and less likely to look at elements in the environment. But because the virtual reflections are references to elements in the machine’s environment, their effects on the distribution of an observer’s attention could be twofold: They could draw attention to the eyes themselves, and they could draw attention to the reflected element in the environment. Mirror Eyes could therefore serve as a tool to establish joint attention between machine agents and human observers. On top of their original role as a general purpose tool for spatial references, an ability to establish and modulate joint attention could extend means for machine-mediated joint attention skill development (e.g., [59]) in people who experience challenges in reading social signals, as seen in individuals on the autism spectrum [60, 61]. Future work should therefore investigate the impact of the Mirror Eyes on people’s attention and whether mirroring can be used for a selective modulation of eye model salience and gaze target salience.